### v1

# Towards Contextual Sensitive Data Detection

Abstract
The emergence of open data portals necessitates more attention to protecting sensitive data before datasets get published and exchanged. While an abundance of methods for suppressing sensitive data exist, the conceptualization of sensitive data and methods to detect it, focus particularly on personal data that, if disclosed, may be harmful or violate privacy. We observe the need for refining and broadening our definitions of sensitive data, and argue that the sensitivity of data depends on its context. Based on this definition, we introduce two mechanisms for contextual sensitive data detection that consider the broader context of a dataset at hand. First, we introduce type contextualization, which first detects the semantic type of particular data values, then considers the overall context of the data values within the dataset or document. Second, we introduce domain contextualization which determines sensitivity of a given dataset in the broader context based on the retrieval of relevant rules from documents that specify data sensitivity (e.g., data topic and geographic origin). Experiments with these mechanisms, assisted by large language models (LLMs), confirm that: 1) type-contextualization significantly reduces the number of false positives for type-based sensitive data detection and reaches a recall of 94% compared to 63% with commercial tools, and 2) domain-contextualization leveraging sensitivity rule retrieval is effective for context-grounded sensitive data detection in non-standard data domains such as humanitarian datasets. Evaluation with humanitarian data experts also reveals that context-grounded LLM explanations provide useful guidance in manual data auditing processes, improving consistency. We open-source mechanisms and annotated datasets for contextual sensitive data detection at https://github.com/trl-lab/sensitive-data-detection.
1 Introduction
Data sharing underpins reuse across sciences, enterprises, and governments through public and private portals [borgman2025data, worth2024datatransparency, zhang2024datamarketssurvey, wilkinson2016fair, brickley2019googledataset]. With the rise of open data sharing portals, comes the need to protect sensitive data from undesired and harmful use. Sensitive data includes direct personal identifiers (e.g., names, emails [gdpr]) as well as indirect or contextual information that becomes a potential risk under certain conditions [unocha2025]. For example, facility locations or military logs may be sensitive depending on geopolitics or timing. This illustrates the complexity of sensitive data detection and the need for nuanced, context-dependent methods. Beyond operational and ethical concerns of publishing sensitive data, Large Language Models (LLMs) increase the risks of leakage: trained on web data, they may memorize and reproduce sensitive information [subramani2023piihuggingface, worth2024datatransparency, carlini2021extracting, lukas2023analyzing]. Studies show sensitive data is often underspecified in documentation on open data portals like HuggingFace [yang2024navigatingdatasetdocumentationsai, croissant2024benjelloun, worth2024datatransparency], illustrating the importance of accurate methods and tools for (semi-)automated sensitive data protection, particularly in open data portals.
Most work on sensitive data detection methods, however, focuses on detecting person-level data, particularly personal identifiable information (PII) [KUZINA2023119924, subramani2023piihuggingface, croissant2024benjelloun]. For unstructured data, PII detection relies on named entity recognition [subramani2023piihuggingface]; similar approaches exist for tabular data [raman2001potter, Hulsebos2019Sherlock], which is the focus in this paper. Yet sensitivity extends far beyond explicit PII. kober2023whatissensitive propose a taxonomy of always-, combination-, context-, and value-sensitive data. For instance, an address may be harmless in the context of a public organizational website but expose a private address in another document, causing naïve type-based detection yield false positives. Likewise, hospital geo-coordinates may be harmless in one region but sensitive in conflict zones [unocha2025], calling for contextualization beyond detecting personal-sensitive data. These limitations call for a more nuanced perspective and framework for assessing sensitive data. LLMs offer new potential for sensitivity detection by reasoning over structured data. Studies show they can interpret tables with carefully designed prompts [fang2024largelanguagemodelsllmstabular], achieving strong results in tasks such as text-to-SQL and table retrieval [fang2024largelanguagemodelsllmstabular, ji2025target]. These capabilities suggest that LLMs can leverage schema, values, and external knowledge to interpret tabular dataset semantics [Hulsebos2019Sherlock].
In this paper, we first revisit the concept of sensitive data, arguing for a more contextual perspective. We introduce two new LLM-assisted mechanisms for contextual sensitive data detection. First, we introduce the detect-then-reflect mechanism for type contextualization of data values within the entire document or tabular dataset to assess their sensitivity. Experiments with real-world datasets containing personal identifiable information show that this mechanism yields a strong improvement in precision while maintaining high recall. Second, the retrieve-then-detect mechanism for domain contextualization which grounds the assessment in domain-specific sensitivity specifications. Evaluations illustrate the potential of this mechanism, achieving high recall and improved consistency for detecting non-personal sensitive data. Importantly, the context-grounded domain-specific explanations were deemed highly useful by experts in humanitarian data assessment in a case-study.
2 Background
Conceptualization of Sensitive Data
The notion of “sensitive data” is multifaceted, spanning legal, theoretical, and domain-specific views. Legal frameworks often equate sensitivity with personal identifiable information. The GDPR defines personal data broadly as “any information relating to an identified or identifiable natural person,” specifying sensitive categories like health, ethnicity, or politics requiring extra protection [gdpr]. Similarly, HIPAA defines identifiers that make health information “Protected Health Information (PHI)” when linked to individuals [hipaa]. These protocols provide a baseline: data that directly or indirectly identifies a person is sensitive. Yet static, type-level definitions are insufficient for modern governance [Quinn_Malgieri_2021], as they emphasize direct identifiers while overlooking non-personal but risky data such as geolocation data, community-level attributes, or institutional records. Sensitivity is often contextual: seemingly harmless data can become sensitive when combined, repurposed, or shared across contexts [malkin2022contextualintegrity].
Theoretical perspectives deepen this nuance. Nissenbaum’s Contextual Integrity frames sensitivity as the appropriateness of information flows within social norms [nissenbaum2004privacy]. Violations occur when flows breach expectations, not from the data type itself. Building on this, researchers advocate context-grounded definitions [malkin2022contextualintegrity, shvartzshnaider2019going]. For instance, shopping data may be ordinary in retail but sensitive in behavioral profiling. Recent taxonomies further refine this view. kober2023whatissensitive distinguish: (1) always-sensitive data (e.g., direct identifiers); (2) combination-sensitive (sensitive when linked together); (3) context-sensitive (sensitivity by use); and (4) value-sensitive (sensitivity from specific values). This shows that not all personal is equally sensitive, and not all sensitive data is on individuals.
In short, conceptualizations have shifted from static identifiers to context-dependent models, where frameworks like Contextual Integrity and multi-layer taxonomies emphasize context, use, and value. This motivates our focus: effective data protection must ensure coverage beyond personal-level information toward context-aware, domain-specific approaches.
Methods for Detecting Sensitive Data
Most automated sensitive data detection methods concentrate on detecting sensitive data of individuals. Widely adopted industry tools like Microsoft Presidio [presidio] and Google Cloud DLP [google_DLP] are largely rule-based, relying on regular expressions, keyword lists, and simple machine learning heuristics to recognize well-defined patterns like email addresses or phone numbers. Research has advanced towards more semantic understanding. Hulsebos2019Sherlock introduced Sherlock, a deep learning model that classifies types of table columns, including numerous PII, by learning from column-level representations. KUZINA2023119924 develop this further with CASSED, a BERT-based model that integrates column headers, data types, and sample values to classify columns into sensitive categories. These approaches outperform regular expressions and dictionary-based baselines, especially on noisy or inconsistently formatted data. Recently, Large Language Models (LLMs) have emerged as powerful tools for sensitive data detection, as particularly shown for PII detection outperforming both rule-based and traditional machine learning methods in accuracy and adaptability[shen2025enhancingdeidentificationpersonallyidentifiable, Yang2023Exploring]. However, a key limitation remains: LLMs excel at identifying known PII types but struggle with assessing contextual sensitivity, such as domain-specific sensitive types, unless explicitly provided with the relevant operational context.
We identify a significant gap in detecting sensitive data: methods focus on detecting personal sensitive data, overlooking non-personal data that is sensitive due to its context. Our work bridges this gap by developing a framework that addresses both personal and non-personal sensitive data detection through contextualization.
3 Revisiting the Concept of Sensitive Data
The main conceptual, legal, and methodological focus regarding sensitive data is on personal- and individual level data, including personal identifiable information (PII) which refers to data that can identify an individual [gdpr]. Regulatory data protocols, such as GDPR [gdpr] and HIPAA [hipaa], specify standards for data sensitivity, hence provide concrete guidelines for data protection measures. However, we argue for the need of a more contextual view of sensitive data (Def. 1) for developing new detection methods.
Definition 1(Contextual Sensitive Data)
Data whose sensitivity depends on external factors and requires protection due to how, by whom, and in what context it can be misused.
Advancing this definition, we posit two contextual aspects that matter in order to assess sensitivity of data: type contextualization and domain contextualization.
Following the existence of generic standards and protocols, sensitive data detection methods often focus on detecting data types reflecting personal sensitive data (e.g., names, addresses, phone numbers, birth dates). But the appearance of a name or address does not immediately make a data value sensitive. For example, an address may reflect the address of a person or organization, each having significantly different consequences, hence different sensitivity levels. We term this type contextualization: whether a data value corresponding to a sensitivity type of interest (e.g. name) truly is sensitive requires considering the context in which it appears in the entire dataset (e.g., the entire document or table).
Besides a more fine-grained contextual notion of a data value its sensitivity, we also argue for a more coarse-grained contextual perspective. Nissenbaum introduced the theory of Contextual Integrity [nissenbaum2004privacy], which states that privacy risks arise when data is used outside its intended setting, particularly addressing personal privacy. This contextual perspective extends beyond personal privacy, in domains such as humanitarian aid, national security, and company-sensitive context. We pose that data protection should account for risks beyond identity exposure, for example, harming vulnerable populations or organizations. We refer to this broader perspective on data sensitivity as domain contextualization: sensitivity emerges from the dataset’s nature and potential misuse.
4 Toward Contextual Sensitive Data Detection
Protecting sensitive data generally follows a three-step procedure: ➀ specifying sensitive data, ➁ detecting sensitive data, ➂ remedying data sensitivity (Figure 1). In this work, we focus on the second step, which has received little attention in research. Existing methods for detecting sensitive data suffer from two key limitations to address the contextual nature of sensitive data:
  1. 1.
    Insufficient specificity: sensitive data detection methods treat all values of types of interest as equally sensitive, while the subject of the value that implies its sensitivity. The lack of specificity leads to low precision hence high false positive rates.
  2. 2.
    Insufficient coverage: sensitive data detection methods typically focus on personal-level data, particularly PII, and do not generalize to domain-specific sensitive data. This lack of coverage leads to low recall of methods, hence leaving sensitive data unprotected.
Figure 1: Data protection process, starting with the specification of sensitivity, detection sensitive data in inputs accordingly, followed by remedying sensitive data by suppression methods.
The contextualization framework introduced in Section 3 addresses these limitations. Type contextualization would improve specificity as potential sensitive values are only considered sensitive based on the broader context of the value. In turn, domain contextualization ensures higher coverage of sensitive data assessment by grounding the assessment in domain-specific context. In this paper, we introduce two concrete mechanisms to address these contextual aspects of sensitive data: the detect-then-reflect mechanism for type contextualization and the retrieve-then-detect mechanism for domain contextualization. In what follows, we specifically focus on tabular data but these mechanisms extend to other modalities, such as free-form text documents.
(a) The detect-then-reflect mechanism for detecting sensitive table columns using type contextualization to increase the precision. This mechanism entails: ➀ detecting potentially sensitive columns based on their type, ➁ having an LLM reflect on the actual sensitivity of the column within the context of the entire table.
(b) The retrieve-then-detect mechanism for detecting sensitive data that requires domain contextualization. This mechanism entails: ➀ retrieving relevant contextual information such as rule-based heuristics, ➁ using an LLM to detect sensitive data using its parametric knowledge and the retrieved expert heuristics.
Figure 2: Mechanisms for detecting sensitive data in a contextual manner.
4.1 Type contextualization: detect-then-reflect
Existing methods typically classify sensitive data based on signals indicating a type of interest, e.g. names, addresses, or phone numbers. If a value matches this type, it is flagged as sensitive. We introduce a two-step pipeline for detecting type-informed sensitive data based on a detect-then-reflect procedure (Figure 2(a)). In the initial detection step, potentially sensitive columns are identified by first classifying columns corresponding to types of interest that are specified upfront, e.g. email, phone number, based on column names and data values. This step prioritizes recall in order to avoid missing any columns that may carry data of the type that can potentially be sensitive. In the second step, reflection, the model assesses the columns in the context of the full table. We prompt an LLM with the table, including the column’s detected type. The LLM determines whether the column is truly sensitive, taking into account inter-column relationships and data across the entire table. This step is crucial for mitigating false positives from the detection stage by incorporating inter-column relationships and table semantics.
4.2 Domain contextualization: retrieve-then-detect
While type-based data sensitivity can often be inferred from structural context, the data sensitivity may depend on domain-specific and geopolitical factors that lie beyond the data itself. For such cases, requiring domain contextualization, we introduce the retrieve-then-detect mechanism (Figure 2(b)). This mechanism considers external contextual information, such as rule-based heuristics specified in data governance documents, in the classification process. Relevant documents are retrieved based on, for example, the origin of a dataset and included in the prompt alongside the table and column under analysis. The LLM leverages both its internal parametric knowledge and the retrieved external data sensitivity instructions to assess whether a column contains data that should be considered sensitive in the given operational context. Unlike type-informed detection, where there is a predefined list of entities (e.g. email, phone number) as relevant for for example PII detection, domain-contextual sensitive data has no fixed ontology. This mechanism establishes domain-specific column sensitivity based on the broader table context and retrieved contextual knowledge.
5 Experimental Setup
Here, we outline the implementation details, the LLMs used, baselines, datasets, and metrics to evaluate our proposed framework for contextually sensitive data detection. Code and datasets can be found at https://github.com/trl-lab/sensitive-data-detection.
5.1 Method implementation
5.1.1 Type-contextualization
In the detection step, given the well-defined nature of PII types, we parse the problem as a multiclass classification task over a taxonomy of personal identifiable types (e.g., email_address, generic_identifier) derived from GDPR [gdpr] and UN OCHA guidelines [OCHA2021Guidelines] (see the code repository for a complete list). The LLM is prompted with a column name and five values, to return only the most appropriate type or None. We consider zero-shot and fine-tuned approaches. The zero-shot method leverages the model’s inherent ability to follow instructions, while the fine-tuned approach enhances performance by training the models on a synthetic dataset of 1,000 labeled columns. Fine-tuning is done using Low-Rank Adaptation [LORA21_HU] for parameter efficiency. The models are optimized with AdamW for a single epoch using an 80/20 train-validation split.
In the reflection phase, the LLM re-evaluates each candidate column corresponding to types of interest within the context of the entire table. We provide the LLM with a markdown-formatted representation of the full table, including headers and five sample rows per column, along with the specific type detected for the target column and, if applicable, the entities of other columns which are detected as PII entities. The LLM is then instructed to assign one of three contextual sensitivity levels: non_sensitive, referring to information that cannot identify a person (e.g., aggregate data, organization names); moderate_sensitive, referring to data that could potentially identify a person when combined with other attributes (e.g., demographics, partial information); and high_sensitive, referring to data that definitively identifies a person (e.g., full name, email, national ID).
5.1.2 Domain-contextualization
For the retrieval step, we aim to identify relevant heuristics indicating sensitivity rules from domain-specific data governance documents in order to integrate this external domain knowledge in the LLM’s context. In our experiments we focus on a case study on the Humanitarian Data Exchange (HDX) [hdx], hence on humanitarian datasets. The system retrieves relevant policy fragments from Information Sharing Protocols (ISPs) based on the dataset’s geographical context (e.g., country). We use GPT-4o-mini to extract the specific sensitivity rules from these documents (e.g., data types for low, moderate, high, severe sensitivity), which we store in a dictionary. If no country-specific ISP is identified, a default fallback template based on cross-country consensus is used.
In the detection phase, the LLM is provided with a curated prompt containing table context (headers and five representative rows), relevant ISP extracts containing explicit instructions to classify each column’s sensitivity according to the ISP guidelines while justifying its decision by directly citing the relevant policy passages. Based on the policy definitions, the model assigns one of four sensitivity levels, non_sensitive, moderate_sensitive, high_sensitive, or severe_sensitive. This mechanism ensures that sensitivity assessments are grounded in explicit domain policies, thereby enhancing both interpretability and practical usability for domain experts.
5.2 Model Selection
We evaluate closed-source and open-source models, to understand trade-offs between these options. From closed-source options, we select the GPT-4o-mini [openai2024gpt4ocard] model for its good performance and cost trade-off. We focus on open-source LLMs that enable local deployment, for which we select the Gemma family (2 and 3) [gemma_2025, Gemma2Paper] and Qwen3 in 8B and 14B sizes [yang2025qwen3technicalreport]. Qwen3 14B was selected given the strong performance of its predecessor Qwen2.5’s exceptional performance on benchmarks such as MMLU Pro  [^1] https://huggingface.co/spaces/TIGER-Lab/MMLU-Pro, demonstrating strong reasoning capabilities and robust general understanding. Furthermore, Gemma2 9B and Gemma3 12B were included to strike a balance between computational efficiency and performance, and their support for quantized deployment. We also include versions of Qwen3 8B and Gemma2 9B that are fFgine-tuned on a synthetic dataset of 1K table columns hand-labeled with 27 PII types, indicated with “FT” suffix.
5.3 Baselines
We compare our methods with two common baselines for sensitive data detection. We include Google Cloud DLP [google_DLP], a rule-based system that combines built-in infoTypes, predefined detectors for common PII entities, which are associated custom-defined dictionaries and regular expressions. For our experiments, we configure the minimum likelihood threshold to “Likely”. The second baseline is Microsoft Presidio, an open-source tool that leverages regular expressions, pattern-based rules, context windows, and ML-based named entity recognition to detect PII types. To enable consistent evaluation, each Presidio type has been mapped to a corresponding category in our unified taxonomy (this mapping can be found in the code repository). Both systems output column-level predictions and are widely adopted in enterprise settings. However, neither tool supports contextual reflection or table-wide reasoning; any match to a known PII pattern is automatically treated as sensitive.
5.4 Datasets
To evaluate the contextual sensitive data detection mechanisms we use a combination of synthetic and real-world datasets, to maintain ground-truth validity for sensitive data detection in generic context and the domain-specific context of humanitarian datasets.
  • Type contextualization with sensitive PII: We extracted 66 anonymized tables (2,061 columns) from the GitTables corpus [hulsebos2023gittables] representing a variety of real-world database schemas. Each column was manually annotated with (i) its PII entity type and (ii) its contextual sensitivity level (non-sensitive, moderate or highly sensitive).
  • Domain contextualization with humanitarian tables: To evaluate the retrieve-then-detect mechanism, we generated 9 synthetic tables with sensitive data closely resembling typical humanitarian data (e.g., for displacement tracking, public health monitoring) for different countries. Each table contains columns that could be potentially sensitive within country-specific Information Sharing Protocols (ISPs). Humanitarian data quality experts labeled each column with a sensitivity level (low, moderate, high, severe). We also add 14 real tables from the Humanitarian Data Exchange [^2] https://data.humdata.org/ that do not contain any sensitive data, to evaluate the false positive rate.
5.5 Metrics
We evaluate along three axes: PII detection, contextual reflection for PII, and non-PII sensitivity detection. To account for class imbalance, we report both weighted (support-weighted) and macro averages. Weighted averages reflect overall performance across examples, while macro averages treat each class equally and reduce dominance by frequent classes. Unless otherwise noted, we focus on weighted averages.
Performance is assessed at both the column and table level. At the column level (type contextualization), PII detection is treated as a multiclass task over 27 PII types (plus a “None” category), while sensitivity reflection reduces to a binary decision between sensitive (moderate/high) and non-sensitive (low). At the table level (domain contextualization), a table is classified as sensitive if it contains any moderate/high/severe sensitive information, resulting in a binary task (sensitive vs. not). Precision, recall, and F1 are used as evaluation metrics.
For labeling, sensitivity levels are binarized: low is treated as non-sensitive, while moderate, high, and severe are grouped as sensitive. This aligns with operational practice where any moderate or higher sensitivity triggers safeguards, consistent with UN OCHA guidance that such data warrants protective measures [unocha2025].
6 Results of contextual sensitive data detection
We present a comprehensive evaluation of our proposed framework for contextually sensitive data detection in tabular datasets, through type- and domain-contextualization. Type Contextualization experiments instrumented by the detect-then-reflect mechanism are focused on personal-identifiable information (PII) and Domain Contextualization instrumented by the retrieve-then-reflect mechanism which focus on less standard data semantics through a humanitarian case-study.
6.1 Type contextualization improves specificity.
First, we assess the detect-then-reflect mechanism for detecting sensitive data through type contextualization. We report on baselines representing existing state-of-the-art tools for detecting sensitive data which automatically treat all columns of interest types as sensitive above the line. Below the line, we report the contextual LLM-assisted method that detects columns of the types of interest which are directly labeled as sensitive without reflection (left column), and are labeled as sensitive only if this is the output of the reflection step (right column).
Table 1: Type contextualization for PII sensitive data detection, with baselines and the detect-then-reflect mechanism. The LLMs are shown with and without the reflection step. We observe that reflection consistently improves precision and F1.
System / Model Without Reflection With Reflection
Prec. Rec. F1 Prec. Rec. F1
Google DLP 0.531 0.628 0.576
Presidio 0.520 0.618 0.565
GPT-4o-mini 0.856 0.639 0.732 0.938 0.632 0.755
Gemma 2 9B 0.740 0.819 0.778 0.800 0.792 0.796
Gemma 3 12B 0.487 0.941 0.642 0.753 0.806 0.779
Qwen3 8B 0.742 0.868 0.800 0.749 0.868 0.804
Qwen3 14B 0.565 0.972 0.714 0.732 0.941 0.824
Qwen3 8B FT $\rightarrow$ GPT-4o-mini 0.902 0.861 0.881
Without reflection, LLMs mark any detected PII-type column as sensitive, leading to high recall but lower precision; Gemma 3 12B and Qwen3 14B reach a recall of 0.94 and 0.97 but precision of only 0.49 and 0.57, respectively. GPT-4o-mini achieves higher precision, 0.86, at lower recall of 0.64. With reflection using full-table context, substantially improves precision across models, e.g., Qwen3 14B increases to 0.73 precision with 0.94 recall, trading a small decrease in recall for a significant gain on precision. A modular pipeline combining Qwen3 8B for type detection with GPT-4o-mini for reflection achieves F1 0.88, but is closely matched by Qwen3 14B which yields higher recall (0.94 compared to 0.86). These results illustrate the stark difference compared to existing baselines reported in the upper left quadrant, which suffer from insufficient recall ánd precision. Comparing the detect-then-reflect mechanism powered by open-source LLMs with commonly used pattern-matching tools DLP and Presidio, we observe a significant improvement in both precision and recall, addressing the key limitations of existing tools for deployment in practice. Overall, the detect-then-reflect mechanism corrects the naive assumption that all columns with potentially sensitive types are sensitive, boosting precision while maintaining recall.
6.1.1 Ablation: reflection without detection.
To investigate the value of the type-detection step, we also task models with identifying sensitive columns using directly the full table context without first detecting the presence of columns of specific types (Table 1). The results in Table 2 illustrate that LLMs yield good performance for sensitive columns using only table context without prior type detection, but generally yield significant improvement across models when explicit types are detected first. Qwen3 14B achieves the highest F1 (0.82) followed by Gemma 2 9B (F1 0.81). GPT-4o-mini excels in precision (0.96) but has low coverage with a recall of 0.56. Comparing the performance with the detect-then-reflect pipeline in Table 1, we observe that the reflection-only approach, without type detection, yields particularly a lower coverage, therefore leaving more sensitive data undetected. For example, looking closer into the performance of the best performing models, the recall of Qwen3 14B drops from 0.94 to 0.89, and Gemma2 9B from 0.79 to 0.75. These results indicate that LLMs can reason about sensitivity from the table alone, but the full detect-then-reflect mechanism yields preferable trade-off in precision and recall.
Table 2: Reflection-only performance without prior type detection step, for identifying contextually sensitive columns. Evaluated across models using precision, recall, and F1 scores.
Model Precision Recall F1
GPT-4o-mini 0.964 0.562 0.711
Gemma 2 9B 0.878 0.750 0.809
Qwen3 8B 0.735 0.781 0.758
Qwen3 14B 0.761 0.885 0.819
6.2 Domain contextualization provides coverage and context-grounded explainability.
Shifting towards contextual sensitive data which cannot be assessed through a fixed set of types, we evaluate the retrieve-then-detect mechanism on a combined set of synthetic and real humanitarian tables from the HDX platform. Without external domain knowledge, i.e. sensitivity heuristics obtained through context retrieval, all models exhibit a highly conservative bias, achieving perfect recall but all with precision lower than 0.56 (Table 3). Integrating retrieved domain-specific context, in this case we leverage information sharing protocols (ISP’s) for humanitarian data sharing, into the prompt consistently improves precision. GPT-4o-mini’s precision increases from 0.47 to 0.69 (F1=0.82), and Qwen3 14B’s from 0.56 to 0.64 (F1=0.78), both maintaining perfect recall (1).
Table 3: Table-level classification performance with and without context-grounding in humanitarian data sensitivity rules. The retrieved rules improve precision while maintaining high recall.
System / Model No Domain Knowledge With Domain Knowledge
Prec. Rec. F1 Prec. Rec. F1
All-tables-sensitive (baseline) 0.375 1.000 0.545
GPT-4o-mini 0.474 1.000 0.643 0.692 1.000 0.818
Gemma 2 9B 0.375 1.000 0.545 0.429 1.000 0.600
Gemma 3 12B 0.529 1.000 0.692 0.500 1.000 0.667
Qwen3 8B 0.562 1.000 0.720 0.778 0.778 0.778
Qwen3 14B 0.562 1.000 0.720 0.643 1.000 0.783
Qualitative analysis reveals that the retrieval mechanism enables models to produce interpretable justifications, thereby enhancing the transparency and auditability of the sensitivity assessments. The below excerpt shows a context-grounded LLM assessment on a table titled “Reported location on NFIs suppliers” with a description of the data at hand (1), along with the LLM’s sensitivity assessment (2), and explanation grounded in retrieved ISP rules (3).
1. The column ‘‘nfi_supplier_loc- value -from_Iran’’ indicates the origin of NFIs suppliers.
2. This information, even with country-level disaggregation, doesn’t reveal personal or sensitive data. It relates to the location of suppliers, which is generally not considered sensitive.
3. It aligns with the ISP’s guideline on ‘‘Facility data (e.g., health, education, water points) at national or regional level, unless explicitly restricted by clusters.’’ Since this column is about locations of suppliers and not about individuals or specific operations, it is categorized as non-sensitive.
The model outputs containing sensitivity classifications and explanations were also reviewed by two humanitarian data quality experts, which considered context-grounded explanations highly informative for consistency and improving human-assessment. Table 3 summarizes results on 9 datasets synthesized based on humanitarian scope, and 14 real humanitarian datasets from HDX without sensitive data to evaluate false positives.
7 Contextual Sensitive Data: importance, practice, and consequences
With the emergence of public data sharing portals, and their adoption in critical areas such as the Humanitarian Data Exchange (HDX) for humanitarian aid coordination, it is key to protect sensitive datasets before they are automatically retrieved from the portal by malicious users or web crawlers for AI training datasets. Not doing so, induces significant risk to vulnerable populations. We reconsider and broaden the definition of what data is sensitive, and argue that sensitive data should be assessed contextually. This shift is critical as prior research for detecting and handling sensitive data typically focuses mainly on personal-level sensitive data. To broaden this practice, we present mechanisms for detecting sensitive data based on the principles of type contextualization and domain contextualization, which form a steppingstone towards this ambition.
On real-world humanitarian datasets, we show that more contextual methods lead to higher precision, hence less false positive alarms, while preserving coverage. These mechanisms can enable automation but also assist humans in data reviewing by providing context-grounded model explanations and concrete citations to relevant sensitivity rules in domain-specific documents. For successful adoption in practice, reasonable sensitivity instructions should be available in the form of 1) fine-grained semantic types of data that are deemed sensitive for type contextualization, and 2) reasonably specified assessment rules in domain-specific documents.
A potential negative consequence of automated data sensitivity detection is that critical datasets might be blocked that are time-sensitive and could be instrumental for solving societal problems, for example, in humanitarian domains. The mechanisms could be tailored around the desired precision/recall trade-off to reduce false alarms. At this moment, the methods are only developed for data sensitivity detection, and not for automated remedies. In practice, a report could be generated based on existing cases to instruct the user on how to remedy the sensitive data or potentially even automatically remedy this.
8 Conclusion
The rise of open data portals are key in facilitating the democratization of access to important data. This development, while positive, comes with risks of having sensitive data being used for harmful purposes by malicious actors as well as indirect leakage into language models which are fueled by any data on the web. Protecting sensitive data from getting published through such portals, by accident, is therefore key. In this paper, we first revisit the concept of sensitive data which is typically focused on personal-level data like personal-identifiable information (PII). Methods for detecting sensitive data are limited: they are too specific to cover all sensitive data, and too generic to only flag data that is actually sensitive. We have introduced a broader, contextual, perspective on sensitive data, and present a framework that formalizes contextual sensitive data from two angles. Experiments show that type-contextualization through an LLM-assisted mechanism that first detects the type of the data values (e.g. names) followed by a reflection on their sensitivity, significantly improves precision and recall over commercial tools and out-of-the-box LLM baselines. A case study with humanitarian datasets on the Humanitarian Data Exchange shows that the domain-contextualization through automated retrieval of rule-based heuristics from domain-specific documents effectively grounds an LLM’s assessment of data sensitivity. With our conceptualization of contextual sensitive data, we intend to encourage researchers and practitioners to consider a broader view on data sensitivity. We believe that our framework for detecting sensitive data provides a stepping stone in this important direction.
Acknowledgements
We are grateful for the collaboration with the Responsible Data team of the United Nations OCHA Humanitarian Data Centre. Specifically, we like to thank Alexandru Gartner, Godfrey Takavarasha, Jos Berens, Javier Téran, Metasebya Sahlu, Melanie Rabier, Nafissatou Pouye, Sarah Telford, Serban Teodorescu, and Shahrooz Badkoubei, for their invaluable contributions to this project. This work has been supported by funding from an AiNed grant from NWO (NGF.1607.22.045), a grant from SAP, and credit grants from OpenAI, Google and Cohere.